mrfunnybunny:

This is for my baby cousin: In 7 days, Lin's will be paid $75 and will deposit it into her checking account. If there are no other transactions besides this deposit and the daily fee, will Lin continue to be charged $7.95 each day after thus deposit is made? What will her balance be at that time?

minustempo:

To make a linear equation, we want the y-intercept (starting deposit) and the slope (daily fee).

minustempo:

final balance = initial deposit - daily fee * number of days

axie:

75-(7.95)*7= Final Balance

axie:

@axie wrote:
75-(7.95)*7= Final Balance
EDITED 75-(7.95*7)= Final Balance

mrfunnybunny:

my cousin: thx